Solve each equation.
-3/x = x/-3

Answer by rfadrogane(214)

Solve each equation.
-3/x = x/-3
Sol'n:
-3/x = x/-3, try to multiply by their LCD: -3x on both side, to eliminate the denominator
-3(-3)= x^2
9=x^2
x=3 ----answer